Logic and Computation

80-310/610, Fall 2001

Course information

Time: Mondays and Wednesdays, 10:30-11:50 Room: Porter Hall 18B Web page: http://www.andrew.cmu.edu/course/80-310-610

Instructor

Name: Prof. Jeremy Avigad Office: Baker Hall 155F email: avigad@cmu.edu Office hours: Monday, 2:30-3:30 and Wednesday, 9-10

Grader

Name: Tianjiao Chu Office: Baker Hall 143 Mailbox: Baker Hall 135 email: tchu@andrew.cmu.edu Office hour: Tuesday, 3–4

Text

Dirk van Dalen, Logic and Structure, and handouts.

Only the van Dalen text is required. But if you are not entirely comfortable reading and writing proofs, take a look at How to Read and Do Proofs by Daniel Solow. If you are not comfortable with first-order logic, you might find Tarski’s World, by Jon Barwise and John Etchemendy, to be helpful.

Requirements

Homework assignments (50%): Due every Wednesday, in class Midterm exam (20%): In class, Wednesday, October 17 Final exam (30%): During the final exam period

Please read the relevant sections of the text before class so that you will have a better understanding of the lecture.

Grading and lateness policy

Homework is due at the beginning of class on Wednesday. It can be turned in late to the grader’s mailbox by noon on Thursday at a 15% penalty, and by noon on Friday at a 30% penalty. Note that this is much better than not turning it in at all, which counts as a score of 0.

The total point score on the homework assignments will vary from week to week. But each homework assignment is ultimately scaled to a score from 0 to 5 (4–5 is an A, 3–4 is a B, etc.) and all count equally towards the final grade. This is done to ensure fairness.

(For example: one week, the total number of points may be 42; and I may announce that 34 and above is an A, 26 and above is a B, 19 and above is a C. In that case, a score of 32 translates to 3 + 6/8 = 3.75, roughly a B+.)

The course is not curved, in that everyone may earn an A, or everyone may fail.

The lowest two homework scores will be dropped. Since you are responsible for all the material on the exams, however, I encourage you to do all the assignments.

Grades are determined based on clarity as well as correctness. You may turn in an answer to a question that, arguably, has all the components of a correct answer; but if the grader has to struggle to understand what you are saying, or read between the lines, or weed out false or irrelevant information, you are unlikely to receive full credit.

You are allowed to work together on the homework assignments, and, in fact, this is encouraged. The only restriction is that when you write up the actual solution you turn in you must do so alone, so that the answer reflects your own understanding. Failure to obey these guidelines constitutes cheating.